# Topological excitations in a spin-orbit coupled spin-1 Bose-Einstein condensate under sinusoidally varying magnetic fields

### v1

@@ 4 Conclusion
In conclusion, the interplay between spin-orbit coupling and an external in-plane magnetic field with sinusoidal variation in the $x-y$ direction in a spin-1 BEC introduces an additional layer of complexity to an already intricate system. The sinusoidal variation of the magnetic field in the $x-y$ plane provides a unique spatial modulation that further influences the ground state of the condensate. The topological excitations in the form of vortex formation occurs at the saddle points of the in-plane magnetic field. Along with the spin-exchange interaction and spin-mixing dynamics, the magnetic field exhibits conjugate vortices and anti-vortices in the $m_{F}=\pm 1$ components and vortex brights in the $m_{F}=0$ component. These density modulations correspond to a wavelength $\lambda$ which is inversely proportional to the non-zero momentum values in the k-space densities.
@@
In this work, we have attempted to showcase numerous cases with various combinations of sinusoidal fields in the $x-y$ direction. We have discussed the ground state structures for all of these choices both with and without the SOC, highlighting the subtle differences in each case. The SOC induces an anisotropy in the system which breaks the rotational symmetry of the system. As a consequence of which, the overall angular momentum is not conserved in the system. Furthermore, the combination of SOC and the spatially varying magnetic field gives rise to spatially modulated spin textures, including the formation of unique skyrmions and anti-skyrmions. These topological defects represent fascinating quantum states within the condensate, showcasing the intricate interplay between the periodicity induced by the sinusoidal magnetic field and the coupling of spin and motion.
@@
The sinusoidal variation introduces periodicity into the system, creating an opportunity to engineer and manipulate the condensate’s properties. This perocidity of magnetic field can be exploited to see the effect of various combinations of in-plane magentic fields. Beyond this, the dynamics of a SO-coupled spin-1 BEC in a polarised sinusoidal magnetic field can be explored. Moreover, it will be interesting to look into the skyrmion dynamics of the SO-coupled system. The ability to tune the sinusoidal variation of the magnetic field in a controlled environment not only provides a deeper understanding of the quantum phases emerging from the interplay but also opens up possibilities for designing novel quantum devices and applications. As investigations into this intricate interplay continue, we anticipate the discovery of new spatially modulated quantum states, offering exciting prospects for both fundamental research and technological advancements.
@@ Acknowledgements
We thank Hari Sadhan Ghosh for many insightful discussions. We acknowledge the National Supercomputing Mission (NSM) for providing computing resources of PARAM Shakti at IIT Kharagpur, which is implemented by C-DAC and supported by the Ministry of Electronics and Information Technology (MeitY) and Department of Science and Technology (DST), Government of India.
Funding information
A.S. gratefully acknowledges the support from the Prime Minister’s Research Fellowship (PMRF), India. S.H. acknowledges the MHRD, Govt. of India for the research fellowship. S.D. acknowledges support from AFOSR FA9550-23-1-0034.

### v2

@@ IV Conclusion
In summary, the interaction between spin-orbit coupling and an external in-plane sinusoidal magnetic field in a spin-1 BEC introduces additional complexity to the system. The periodic modulation of the magnetic field in the $x$-$y$ plane creates a distinct spatial variation that influences the condensate’s ground state. Vortex formations occur at saddle-points of the in-plane magnetic field, adding to the system’s topological excitations. Coupled with spin-exchange interactions and spin-mixing dynamics, the magnetic field induces conjugate vortices and anti-vortices in the $m_{F}=\pm 1$ components, and vortex brights in the $m_{F}=0$ component. These density modulations correspond to a wavelength $\lambda$ inversely proportional to the non-zero momentum values at which additional density peaks emerge in the k-space.
@@
In the presence of SOC, the rotational symmetry of the system breaks giving rise to certain topological excitations in the system. Due to the symmetry-breaking, an anisotropy is induced in the system which renders a very unique density distribution to three components of the spin-1 BEC. As a consequence, the overall angular momentum is not conserved in the system. Furthermore, the combination of SOC and the spatially varying magnetic field gives rise to spatially modulated spin textures, including the formation of unique configuration of skyrmions and anti-skyrmions. These topological defects represent fascinating quantum states within the condensate, showcasing the intricate interplay between the periodicity induced by the sinusoidal magnetic field and the coupling of spin and motion.
@@
The sinusoidal variation introduces periodicity into the system, creating an opportunity to engineer and manipulate the condensate’s structure and properties. Beyond this, the dynamics of a SO-coupled spin-1 BEC in a polarised sinusoidal magnetic field can be explored. Moreover, it will be interesting to look into the skyrmion dynamics of the SO-coupled system. The ability to tune the sinusoidal variation of the magnetic field in a controlled environment not only provides a deeper understanding of the quantum phases emerging from the interplay but also opens up possibilities for designing novel quantum devices (e.g., skyrmion qubits [108]) and applications.
Understanding the interplay between SOC and sinusoidal magnetic fields in spin-1 BECs not only advances fundamental physics but also paves the way for new quantum technologies. By investigating these systems, researchers can develop more sophisticated methods for manipulating quantum states, explore the stability and dynamics of topologically non-trivial excitations, and potentially discover new quantum phases with practical applications. Our findings suggests ways to engineer diverse topological excitations in a spin-1 BEC employing a periodically varying magnetic field. This may serve as an important benchmark for future theoretical and experimental studies aimed at harnessing quantum mechanics for technological innovation.
@@ Acknowledgements
We thank Hari Sadhan Ghosh for many insightful discussions. We acknowledge the National Supercomputing Mission (NSM) for providing computing resources of PARAM Shakti at IIT Kharagpur, which is implemented by C-DAC and supported by the Ministry of Electronics and Information Technology (MeitY) and Department of Science and Technology (DST), Government of India. AS gratefully acknowledges the support from the Prime Minister’s Research Fellowship (PMRF), India. SH acknowledges the MHRD, Govt. of India for the research fellowship. SD acknowledges support from AFOSR FA9550-23-1-0034.
Appendix A Impact of varying spin-orbit coupling strength on the spin texture
In section III of the main text, we discussed the interplay of the SOC and magnetic field on the spin-1 BEC for a fixed SOC strength $\gamma=0.5$ for various cases of gradient and sinusoidally varying in-plane magnetic field. Here, we show the impact of variation of the SOC strength on the spin texture. Fig. 15 displays the spatial distribution of the spin texture under the influence of the polarized gradient magnetic field (as discussed in III.1) with varying SOC strengths. The polarised gradient magnetic field applied along $x$-axis ($y$-axis) renders skyrmions (anti-skyrmions) along the $x=0$ ($y=0$) line in presence of the SOC. The number of such structures increases while the wavelength of density modulation decreases with increasing SOC strength, $\gamma$.
Figure 15: (a1-a3) shows the spin texture for the $x$-polarised gradient magnetic field for varying SOC strength $\gamma=1.00,1.50,2.00$ respectively. (b1-b3) shows the same for the $y$-polarised gradient magnetic field. Units are dimensionless.
Figure 16: The spin texture with various SOC strengths $\gamma$ for the cases of sinusoidal magnetic fields discussed in section III.2. The columns present the case specific results while the rows represent different $\gamma$ values. Units are dimensionless.
For the cross-polarised sinusoidally varying in-plane magnetic fields discussed in section III.2, we present the spin textures for varying SOC strengths in Fig. 16. We observe that for different SOC strengths, the spin texture follows a similar trend in both $x$ and $y$-directions for each case. However, the density modulations increase with the $\gamma$ values. As a result of the grwoing anisotropy due to increased SOC strength, the system exhibits a complicated spin texture with more number of skyrmions and anti-skyrmions. Moreover, the interaction between these skyrmions and anti-skyrmions affect their arrangement giving rise to a complex eolotropic structure. As investigations into this intricate interplay continue, we anticipate the discovery of new spatially modulated quantum states, offering exciting prospects for both fundamental research and technological advancements.